# Global Hash Tables Strike Back! An Analysis of Parallel GROUP BY Aggregation

### v1

@@ 3.1. Ticketing
Figure 4. Scaling behavior of various hash maps for ticketing across different data distributions (see Section 4.1 for workload details). The top row measures performance as throughput (higher is better) against thread count. The bottom row plots the speedup factor against thread count, measured as the single-threaded latency divided by the latency at a given thread count. Ideal scaling is the linear function $speedup=threads$, which is plotted as a dashed line. The method with the most ideal speedup factor varies by workload, but Folklore* consistently achieves the best overall throughput across the tested range of thread counts.
@@
In Figure 4, we show the performance and scaling behavior of these designs across different cardinalities and data distributions (see Section 4.1). We find that Folklore* consistently performs best across all tested hash tables, despite its simple implementation. It also exhibits excellent scaling behavior for low cardinality workloads, achieving a 25.6x speedup at 32 threads. Folklore*’s scaling is not ideal at higher cardinalities but still achieves non-trivial speedups (10.6x for high cardinality and 6.0x with unique keys). Further, its efficient implementation still makes it an ideal choice at even high thread counts despite its lower speedup factors. Crucially, the performance of the table is resilient to data skew due to its fast path lookup–in fact, Folklore* performs even better in the presence of heavy hitters due to greater cache locality.
@@
Surprisingly, cuckoo hashing exhibits very poor scaling in high contention workloads, despite us implementing a fast path lookup and prior literature indicating good performance (Li et al., 2014). We attribute this behavior to the implementation using fine-grained locking on buckets. Even though we only acquire read locks on lookups, the bookkeeping required by the locks still creates significant contention when multiple threads access the same resource in a short period of time. The implementation of DashMap also uses read locks which could explain its similar performance characteristics. In contrast, LeapMap and Folklore* are fully lock-free (and Iceberg is lock-free on its fast path lookup). These characteristics explain the significant divide in scaling behavior in the low cardinality case, with the speedup factor at high thread count being starkly higher for the latter three methods compared to the former two that use read locks. Thus, it is essential to ticketing performance to have a “fast path” for reading previously-inserted values, and preferably, this “fast path” should not require taking any locks.
@@ Discussion
An interesting finding from the experiment results is that for our lookup and insert-only workload, even a very simple linear probing design (Folklore*) achieves excellent performance. A contributing factor to this surprising fact is that the typical downside to linear probing, deletions, is a non-issue given our workload. Thus, linear probing’s cache-conscious forward scan does not really have a downside. We conclude that to perform efficient ticketing, fancy hash tables are not required: linear probing is all you need.
That being said, there is likely room for more complicated designs to optimize lookup and insert operations. Unfortunately, evaluating existing state-of-the-art hash table designs and their suitability for ticketing is complicated by gaps in benchmarking results. Existing literature on concurrent hash table designs often does not test certain workloads important for our application. For example, neither of (Li et al., 2014; Kelly et al., 2020) test a lookup-only workload (aligning with low cardinality workloads), and all workloads tested contained updates and deletions (of which there are none in group aggregation). While this particular selection of workloads makes sense for a general-purpose hash table, the results do not apply to group aggregation. Put another way, the omission of read-heavy, delete-free workloads can cause the designs to be poorly optimized for our ticketing use case despite strong performance in mixed, general-purpose workloads.
@@
Finally, we make the observation that our definition of ticketing reduces the task to that of finding a perfect hash function; that is, the purpose of ticketing is to assign each key a unique, (near-)densly packed integer, which is exactly what a perfect hash function does. Gaffney and Patel (Gaffney and Patel, 2024) found that significant speedups can be achieved when integrating perfect hash functions into DuckDB’s aggregation pipeline. Our formulation of ticketing aligns perfectly with this notion of perfect hashing, and ticketing could likely also greatly benefit from perfect hashing (i.e., “skipping” the ticketing step). Although building a perfect hash function requires knowing all the data ahead of time, which violates the morsel-driven parallelism model, it could be highly beneficial in situations where building a computationally efficient perfect hash function is worth the effort (perhaps if queries are frequent on a fixed set of keys). Perfect hashing could essentially remove all contention from the ticketing phase, greatly accelerating fully concurrent aggregation.
@@ 3.2. Partial Aggregate Update
In the update step, described in Section 2, we “actually do” the aggregation and modify the partial aggregate value corresponding to each ticket based on the associated row being aggregated. Whereas in the ticketing step, we did not need to perform any concurrent updates, now concurrent updates are necessary for each incoming tuple. That is, concurrency control problems that we avoided in the ticketing step have been deferred to this stage. In this section, we explore two classes of update methods, concurrent and thread-local. We evaluate and discuss the methods and identify the situations where each exhibits good performance and scaling.
@@
Another simple approach is to have each partial aggregate be an atomic. It is straightforward to perform aggregation functions such as COUNT, SUM, and MIN/MAX, but other more complicated aggregation functions may not be easy to implement with atomics. One general solution is to perform a lookup on the current partial aggregate, perform the update, and then use a CAS instruction to update the partial aggregate. This method, however, increases the number of atomic operations and hence contention, as well as potentially suffering from the ABA problem (Dechev et al., 2010) (although this is probably unlikely for most aggregation functions), so some care is required in implementation. These fully concurrent update methods are simple to implement and very memory efficient, however, they suffer at higher levels of contention: if there is a heavy hitter in the data, every thread may simultaneously and repeatedly try to update the same partial aggregate value, presenting an issue for skewed workloads.
@@
To mitigate issues with contention, we explore a thread-local approach to updates where each worker thread updates partial aggregates in its own thread local vector, eliminating all contention effects in the update stage at the cost of a merge of all partial aggregate values at the end of the aggregation. Although the total work of the merge step scales with the number of threads, because the vectors of partial aggregates are all in the same order (ordered by ticket), the merge is trivially parallel and cache efficient, which should in part mitigate the overhead. This merge can be viewed as the “transpose” of the merge done at the end of the partitioned case: in the partition case, each worker has fully aggregated values but only some of the keys, whereas in the fully concurrent case, each worker has partially aggregated values but all of the keys. Unfortunately, the work per thread does not decrease asymptotically as threads increase. Since each of $k$ threads is assigned $n/k$ tickets to merge (where $n$ is the number of unique keys), and there are $k$ threads worth of partial aggregates per ticket, each thread aggregates over $(n/k)*k$ partial aggregates, yielding $O(n)$ runtime. The runtime per thread of the merge step is constant per thread, and thus merging does not scale at all.
Another challenge with thread local aggregation is that memory usage scales linearly with the number of threads and distinct keys, requiring a vector the size of the entire key space per thread. For very large datasets at high thread counts, this overhead could be a concern, especially since there is no clear way to spill non-partitioned hash aggregations efficiently to disk. One mitigating factor is the fact that vectors of partial aggregates are dense, as opposed to the hash tables allocated per thread in partitioned aggregation, reducing the gap between the two techniques’ memory usage. We quantify the memory overhead in Section 4.5.
@@
Figure 5. Scaling behavior of various partial aggregate update methods across different data distributions. Similarly to Figure 4, the top row plots performance as throughput and the bottom row plots the speedup factor relative to single-threaded performance. Atomic updates display good performance and scaling in low contention scenarios while thread local updates have good performance and scaling for lower cardinality workloads. Thread local updates are also resilient to data skew.
@@
In Figure 5, we plot the performance of these update methods in isolation–that is, without ticketing. Keys are given as an integer from 0 to the max key, which is used directly as a ticket (i.e., a perfect hash function). We set up the experiment in this way to isolate the scaling behavior of each specific aggregation method. It is important to note a caveat that in actuality, contention effects are far less pronounced than how they appear in these isolated benchmarks because the time it takes to ticket causes fewer threads to be at the update step at the same time, decreasing contention.
@@
Both thread local and atomic updates demonstrate strong scaling behavior in some workloads but neither is a clear winner across all workloads. Thread local updates have superior performance in the presence of high contention (i.e. low cardinality datasets or skewed distributions) but degrade in performance as cardinality increases. This is attributable to the fact that at lower numbers of unique keys, the underlying vectors of partial aggregates are small enough that there is minimal overhead from merging. At low cardinality, thread local updates achieve a substantial speedup of 12.6x at 32 threads.
@@
At higher cardinalities, the thread local method demonstrates anti-scaling behavior at higher thread counts, actually decreasing in performance after either 8 or 16 threads. This effect is especially pronounced in the unique key workload, so much so that there is close to no speedup once reaching 32 threads. The performance in the high cardinality workload is more mixed, with thread local updates obtaining a 4.4x speedup at 32 threads but also demonstrating anti-scaling behavior past 16 threads. Crucially, though, the speedup is consistent no matter the skew–since the partial aggregates have no contention, skew can only help due to greater cache locality and leads to no performance degradation. This property makes thread local updates a good option even with relatively high cardinality workloads due to its consistency.
@@
Meanwhile, atomic updates exhibit better scaling in workloads with low contention and high cardinality. For the high cardinality and unique key workloads, when in the absence of skew, atomic updates are a clear winner. In fact, in the high cardinality case, atomics achieves a blazing 20.5x speedup at 32 threads. Surprisingly, atomic updates exhibit good scaling even in the presence of some skew (the Zipfian workload) but fall short when the skew becomes too great (the heavy hitter workloads).
Fine-grained locked updates exhibits similar behavior to atomic updates (since it is bottle-necked by similar contention effects) but tends to be the worst performer due to its higher overhead. However, fine-grained locking exhibits better scaling behavior (but not better absolute performance on today’s hardware) than thread local updates at high thread count and has the added benefit of being able to easily support arbitrary update functions. As hardware continues to add more cores, fine-grained locking is a promising technique for scalability, although other techniques have superior absolute performance on today’s chips.
@@
Table 1. Desiderata fulfilled by each different update method. For cardinality and skew, the range of workloads where the given update method performs well in is given. For memory usage, the asymptotic behavior is given as a factor of $n$ (the number of unique keys) and $k$ (the number of threads).
@@
Discussion
Given the varied performance characteristics of these update methods, summarized in Table 1, ideal performance can best be achieved by adaptively choosing the best method for a given workload. Although imperfect, database optimizers have grown sophisticated, with some work showing good results in predicting sophisticated statistics including the number of unique keys and data skew (Kipf et al., 2019) to efficiently optimize group by aggregations. This work can enable the optimizer to choose the update method on a per-query basis. Absent an adaptive choice, thread local aggregation provides good performance on today’s hardware in all cases but when keys are strictly unique.
There is a concerning performance gap in the case of high skew and high cardinality datasets, which is bad for all methods. The issue is in part mitigated by the fact that the higher the proportion of unique keys, necessarily the lower the proportion of keys that could be duplicated and cause contention. Furthermore, it is important to note poor scaling in the partial aggregate update step does not preclude overall good performance from fully concurrent aggregation. In Section 4.3, we find that the ticketing stage typically accounts for the bulk of overall execution time (since hash table operations tend to have much higher overhead than updates on a vector). Therefore, since we have found that ticketing exhibits good scaling behavior, provided that the anti-scaling in the update step is not bad enough to be a bottleneck, the overall system still can scale well even with degraded performance in the update step.
We believe there is significant room for future work to improve this step. Particularly, a system that can combine both atomic or locked updates with thread local updates could take advantage of the benefits of both systems. This discussion bears similarity to hybrid aggregation approaches (Cieslewicz and Ross, 2007; Fent and Neumann, 2021) that combine a shared hash table with thread local hash tables for heavy hitters. We also believe there can be room for optimization by taking advantage of the vectorized nature of these updates. If locks can be obtained a vector at a time, we may be able to reduce locking overhead. Unfortunately, there appears to be little literature that explores vectorized concurrent updates on an array of values.
@@ 4. End-to-end experiments
In this section, we combine the ticketing and update steps of fully concurrent aggregation and benchmark it end-to-end, comparing performance to an implementation of partitioned aggregation.
Our experimental analysis is broken into five sections. Setup (4.1) details the dataset and workload characteristics we benchmark. Scaling (4.2) analyzes the results of experiments measuring the throughput and scaling behavior of various aggregation methods across different workloads. Latency Breakdown (4.3) aims to explain the observed scaling behavior of fully concurrent aggregations by breaking down the amount of time spent in each stage of the aggregation process. Resizing (4.4) compares aggregation methods in the case of poor cardinality estimation that results in resizing behavior. Memory Usage (4.5) compares the peak memory usage of each aggregation method.
4.1. Setup
We run our experiments on a machine with 256 GB of RAM and a AMD EPYC 9454P processor with 48 cores @ 2.75GHz. All experiments were implemented in Rust and were compiled in release mode using the target-cpu=native flag.
Datasets
Our tests use synthetic datasets consisting of 100 million keys. We varied the cardinality of the dataset, labeled low, high, and unique which consist of 1000 unique keys ($\sim$100lookup/$\sim$0insertion on ticketing), 10 million unique keys ($\sim$90lookup/$\sim$10insertion on ticketing), and 100 million unique keys ($\sim$0lookup/$\sim$100insertion) respectively. On the high cardinality dataset, we also add two types of skew: a Zipfian distribution with exponent parameter $s=0.8$, as well as a heavy-hitter dataset where 50% of the dataset consists of the same key.
Workload
As a demonstrative example of common aggregation functions, we use COUNT for all tests. All experiments (except otherwise noted) assume perfect cardinality estimation and thus perfectly size hash tables and partial aggregate vectors. We later evaluate the effect of inaccurate cardinality estimation and resizing in Section 4.4. For all fully concurrent aggregations, we use Folklore* for the ticketing step because it achieved the highest throughput across all tested workloads. All results (including those given in earlier sections) are obtained by taking the median latency of a given workload after 9 runs (not including warm-up runs).

### v2

@@ 3.1. Ticketing
Scaling behavior of various hash maps for ticketing across different data distributions (see Section 2.4 for workload details). The top row measures performance as throughput (higher is better) against thread count. The bottom row plots the speedup factor against thread count, measured as the single-threaded latency divided by the latency at a given thread count. Ideal scaling is the linear function $speedup=threads$, which is plotted as a dashed line. The vertical dashed line represents the threshold thread count between physical cores and SMT. Folklore* consistently achieves the best overall throughput across the tested range of thread counts and displays good scaling behavior.
Figure 4. Scaling behavior of various hash maps for ticketing. The top row measures performance as throughput (higher is better) against thread count. The bottom row plots the speedup factor against thread count, measured as the single-threaded latency divided by the latency at a given thread count. Ideal scaling is the linear function $speedup=threads$, which is plotted as a dashed line. The vertical dashed line represents the threshold thread count between physical cores and SMT.
Ticketing Performance
@@
In Figure 4, we show the performance and scaling behavior of these designs across different cardinalities and data distributions (see Section 2.4). We find that Folklore* consistently performs best across all tested designs, despite its simple implementation. Crucially, Folklore* performance is resilient to data skew due to its fast path lookup. It also exhibits excellent scaling behavior for low cardinality workloads, although the speedup it achieves at high thread counts degrades as cardinality increases. At 48 threads (the number of real cores on the system), it achieves a 37.6x speedup for low cardinality, 22.8x for high cardinality and 13.8x with unique keys. SMT generally yields little benefit.
@@
Surprisingly, cuckoo hashing exhibits very poor scaling in high contention lookup workloads, despite us implementing a fast path and prior literature indicating good performance (Li et al., 2014). We attribute this behavior to the use of fine-grained locks on buckets. Even though read locks are used on the fast path, bookkeeping by the locks may cause contention when multiple threads access the same resource in rapid succession. DashMap also uses read locks, which could explain its similar performance characteristics.
[…]
Breakdown of time spent on work by processor per workload and thread count. Shows both idle time and topdown analysis of time performing work. Shows the amount of time spent on work and the performance characteristics of the work. See Section 4.2 for more details on this mode of analysis.
Figure 5. Breakdown of time spent on work by processor per workload and thread count. Shows both idle time and topdown analysis of time performing work (see Section 4.2).
In Figure 5, we perform topdown analysis (Yasin, 2014) on a selection of the hash table designs to understand the performance characteristics of each method. More detailed explanation of our methodology can be found in Section 4.2. Our hypothesis regarding bookkeeping on locks aligns with the spike in backend bound behavior when multi-threading for cuckoo hashing on the low cardinality dataset (contention on atomics, used for bookkeeping in read locks, is considered backend bound). Similar behavior occurs for the heavy hitter dataset and for DashMap (not shown).
In contrast, LeapMap and Folklore* are fully lock-free (and Iceberg is lock-free on its fast path lookup). These characteristics explain the gulf in scaling behavior in the low cardinality case, where speedup at high thread count is starkly higher for the latter three methods compared to the former two that use read locks. Further, by avoiding write locks with all methods, we avoid idle time from contention. Thus, it is essential to have a lockless “fast path” for reading previously-inserted values.
Figure 5 also reveals that Folklore* shows less backend bound behavior compared to other methods, suggesting superior cache performance, a benefit of linear probing. In fact, when looking at raw performance counters, across the data distributions and thread counts (up to 48) in all but two cases Folklore* achieves the lowest number of cache misses, and in the majority has the lowest cache miss rate. For example, for the high cardinality dataset at 48 threads, Folklore* has 1.4e8 cache misses and a 39.9% cache miss rate. The next best by total cache misses (Leapfrog) has considerably higher cache miss count and miss rate at 2.1e8 and 46.2% respectively. That said, Folklore* still spends a substantial proportion of time backend bound at high cardinalities, explaining the poorer scaling as cardinality increases, as memory now becomes a bottleneck.
@@ Discussion
An interesting finding from the experiment results is that for our lookup and insert-only workload, even a simple linear probing design (Folklore*) achieves excellent performance. A contributing factor is that the typical downside to linear probing, deletions, is a non-issue given our workload. We conclude that to perform efficient ticketing, fancy hash tables are not required: linear probing is all you need.
One meta-explanation for Folklore*’s dominance is the benchmark workloads used by hash table researchers. Existing literature on concurrent hash table designs often do not test certain workloads important for our application. For example, neither of (Li et al., 2014; Kelly et al., 2020) test a lookup-only workload (aligning with low cardinality workloads), and all workloads tested contained updates and deletions (of which there are none in our formulation of group aggregation). While this particular selection of workloads makes sense for a general-purpose hash table, the results do not apply to group aggregation. Put another way, the omission of read-heavy, delete-free workloads can cause the designs to be poorly optimized for our ticketing use case despite strong performance in mixed, general-purpose workloads.
@@
Finally, we make the observation that our definition of ticketing reduces the task to that of finding a perfect hash function; that is, the purpose of ticketing is to assign each key a unique, densely packed integer, which is exactly what a perfect hash function does. Gaffney and Patel (Gaffney and Patel, 2024) found that significant speedups can be achieved when integrating perfect hash functions into DuckDB’s aggregation pipeline. Our formulation of ticketing aligns perfectly with this notion of perfect hashing, and ticketing could likely also greatly benefit from perfect hashing (i.e., “skipping” the ticketing step). Perfect hashing could essentially remove all contention from the ticketing phase, greatly accelerating fully concurrent aggregation.
@@ 3.2. Partial Aggregate Update
In the update step, we “actually do” the aggregation and modify the partial aggregate value corresponding to each ticket based on the associated row. Concurrency control problems that we avoided in the ticketing step have been deferred to this stage. In this section, we explore two classes of update methods, concurrent and thread local. We evaluate and discuss the methods and identify the situations where each exhibits good performance and scaling.
@@
Another simple approach is to have each partial aggregate be an atomic. It is straightforward to perform aggregation functions such as COUNT, SUM, and MIN/MAX, but other more complicated aggregation functions may not be easy to implement with atomics. One general solution is to perform a lookup on the current partial aggregate, perform the update, and then use a CAS instruction to update the partial aggregate. However, this method increases the number of atomic operations, and hence contention, and also potentially suffers from the ABA problem (Dechev et al., 2010) (although this is probably unlikely for most aggregation functions). These fully concurrent update methods are simple to implement and very memory efficient, however, they suffer at higher levels of contention: if there is a heavy hitter in the data, every thread may simultaneously and repeatedly try to update the same partial aggregate value. There is potential for operation-specific optimizations, such as the inclusion of a fast-path for MIN/MAX that avoids an atomic write if not needed, but in the general case skewed workloads continue to present issues.
@@
To mitigate issues with contention, we explore a thread local approach to updates where each worker thread updates partial aggregates in its own thread local vector, eliminating all contention at the cost of a merge of all partial aggregate values at the end of the aggregation. Although the total work of the merge step scales with the number of threads, because the vectors of partial aggregates are all in the same order (ordered by ticket), the merge is trivially parallel and cache-efficient, which should in part mitigate the overhead. Unfortunately, the work per thread does not decrease asymptotically as threads increase. Since each of $k$ threads is assigned $n/k$ tickets to merge (where $n$ is the number of unique keys), and there are $k$ threads worth of partial aggregates per ticket, each thread aggregates over (n/k)k(n/k)*k partial aggregates, yielding $O(n)$ runtime.
Another challenge is that memory usage scales linearly with the number of threads and distinct keys, requiring a vector the size of the entire key space per thread. For very large datasets at high thread counts, this overhead could be a concern. One mitigating factor is the fact that vectors of partial aggregates are dense, as opposed to the hash tables allocated per thread in partitioned aggregation, reducing the gap between the two techniques’ memory usage. We quantify the memory overhead in Section 4.6.
@@
Scaling behavior of various partial aggregate update methods across different data distributions. Similarly to Figure 4, the top row plots performance as throughput and the bottom row plots the speedup factor relative to single-threaded performance. Atomic updates display good performance and scaling in low contention scenarios while thread local updates have good performance and scaling for lower cardinality workloads. Thread local updates are also resilient to data skew.
Figure 6. Scaling behavior of various partial aggregate update methods across different data distributions.
@@
In Figure 6, we plot the performance of these update methods in isolation–that is, without ticketing. Keys are given as an integer from 0 to the max key, which is used directly as a ticket (i.e., a perfect hash function). We set up the experiment in this way to isolate the scaling behavior of each specific aggregation method. It is important to note that contention effects are far less pronounced than how they appear in these isolated benchmarks because the time it takes to ticket causes fewer threads to be at the update step at the same time, decreasing contention.
@@
Both thread local and atomic updates demonstrate strong scaling behavior in some workloads but neither is a clear winner across all workloads. Thread local updates have superior performance in the presence of high contention (i.e., low cardinality datasets or skewed distributions) but degrade in performance as cardinality increases. This is because, at lower numbers of unique keys, the underlying vectors of partial aggregates are small enough that there is minimal overhead from merging. At low cardinality, thread local updates achieve a substantial speedup of 33.7x at 48 threads.
@@
At higher cardinalities, the thread local method demonstrates inverse scaling behavior at higher thread counts, actually decreasing in throughput after 32 threads for the high cardinality case and 8 threads in unique key case. Still, at its peak, the thread local achieves a 9.2x speedup (at 32 threads) on the high cardinality case. The unique case has a peak speedup of only 3.5x (at 8 threads). Crucially, though, the speedup is consistent no matter the skew–since the partial aggregates have no contention, skew can only help due to greater cache locality and leads to no performance degradation. This property makes thread local updates a good option even with relatively high cardinality workloads due to its consistency.
@@
Meanwhile, atomic updates exhibit better scaling in workloads with low contention and high cardinality. For the high cardinality and unique key workloads, when in the absence of skew, atomic updates are a clear winner. In fact, in the high cardinality case, atomics achieve a blazing 34.7x speedup at 48 threads. Surprisingly, atomic updates exhibit serviceable scaling even in the presence of some skew (achieving 10.3x speedup on the Zipfian workload) but fall short when the skew becomes too great (the heavy hitter workload). Fine-grained locked updates exhibits similar behavior to atomic updates (since it is bottlenecked by similar contention effects) but tends to be the worst performer due to its higher overhead.
Discussion
@@
Table 2. Desiderata fulfilled by each update method. For cardinality, skew, and thread count, the range of workloads where the given update method performs well in is given. For memory usage, the asymptotic behavior is given as a factor of $n$ (the number of unique keys) and $k$ (the number of threads).
@@
Given the varied performance characteristics of these update methods, summarized in Table 2, ideal performance can be achieved by adaptively choosing the best method. Although imperfect, optimizers can estimate the number of unique keys (Kipf et al., 2019) to choose the update method on a per-query basis.
There is a concerning performance gap for workloads that are both high cardinality and highly skewed. However, we note that the most extreme performance degradation occurs in rarer cases, such as when there are a particularly high number of duplicate values, or when elements are unique. Furthermore, it is important to note poor scaling in the partial aggregate update step does not preclude overall good performance from fully concurrent aggregation. In Section 4.2, we find that the ticketing stage typically accounts for the bulk of overall execution time (since hash table operations tend to have much higher overhead than updates on a vector). Therefore, since we have found that ticketing exhibits good scaling behavior, provided that the poor scaling in the update step is not bad enough to be a bottleneck, the overall system still can scale well even with degraded performance in the update step.
We believe that there is significant room for future work to improve this step. Particularly, a system that can combine both atomic or locked updates with thread local updates could take advantage of the benefits of both systems. This discussion bears similarity to hybrid aggregation approaches (Cieslewicz and Ross, 2007; Fent and Neumann, 2021) that combine a shared hash table with thread local hash tables for heavy hitters. We also believe that there can be room for optimization by taking advantage of the vectorized nature of these updates. If locks can be obtained a vector at a time, we may be able to reduce the overhead from locking.
@@ 4. End-to-end experiments
In this section, we combine the ticketing and update steps of fully concurrent aggregation and benchmark end-to-end performance. We also implement partitioned aggregation in the same environment and compare the results. For all fully concurrent aggregations, we use Folklore* for the ticketing step because it had the highest throughput across all tested cases.
Recall that our experimental setup is described in Section 2.4. Our experimental analysis is broken into four sections. Scaling (4.1) presents experiments measuring throughput and scaling behavior across different workloads. Explaining Scaling (4.2) aims to explain the observed scaling behavior by analyzing time spent on each stage of aggregation and profiling results. Other Platforms (4.3) replicates the main experiment on other architectures and discusses any differences in performance. Tuple Size (4.4) varies the size of the aggregate tuple. Resizing (4.5) compares aggregation methods in the case of poor cardinality estimation resulting in resizing behavior. Memory Usage (4.6) compares the peak memory usage of each aggregation method.